Suppose we have a market demand Q = 18 – P and a cost C(Q) 9) = 3Q?.1 Suppose we realize that the market described in question 1 (Market demand is still Q = 18 – P) has a negative externality.
Answer #1

Marginal social benefit is P = 18 - Q. Marginal private cost = Q and marginal external cost = 2Q. This indicates that marginal social cost = Q + 2Q = 3Q

A) the marginal cost of externality is the derivative of the cost of externality function dCE/dQ = 2Q.

B) marginal cost to the society is the sum of private marginal cost and marginal external cost = Q + 2Q = 3Q

C) socially optimal quantity is the one at which marginal social cost and marginal social benefit are equal to each other

18 - Q = 3Q

Q = 4.5 units

P = 13.5 per unit